At its present level of output of 100 units, a perfectly competitive firm discovers that (i) its total fixed costs are $200 and (ii) its marginal cost is $7 and equal to average total cost. At an output level of 50 units, marginal cost is $4 and equal to average variable cost. The price of the commodity being produced is $6.

At present level of output, the firm experiences

A) losses equal to its total fixed cost.

B) zero profits.

C) losses less than its total fixed cost.

D) losses greater than its total fixed cost.
Answer #1

The given information is

Q= 100 units,

a perfectly competitive firm discovers that

TFC= $200 and

marginal cost is $7

average total cost=$7

TC=ATC*Q

=7*100

=700

TR=P*Q

=6*100

=600

Economic Profit=TR-TC

=700-600

=$100

At an output level of 50 units,

marginal cost is =$4

average variable cost=$4

The price of the commodity being produced is $6.

TFC=200

TVC=AVC*Q

=4*50

=200

TR=6*50

=300

Loss= TC-TR

=200+200-300

=-$100

TFC=$200

Hence it can be said that at present level of output, the firm experiences losses less than its total fixed cost.

Hence option C is the correct answer.